# Multi-Crit: Benchmarking Multimodal Judges on Pluralistic Criteria-Following

### v1

@@ D.2 Full Evaluation Instances
Response B
(yi-vl-plus)
In a vividly colored comic book panel, we see two characters engaging in a lively conversation. One character, with a bright yellow head and a red shirt, exclaims, “An inventor? I don’t believe you!” His name, Marco, is displayed in a white rectangle. Opposite him, another character, sporting a black hairstyle and a blue shirt, confidently asks, “Why not?” His name, Babi, is also shown in a white rectangle. The contrasting colors of their clothes and hair add to the dynamic energy of their exchange, as they stand against a plain blue backdrop. The speech bubbles capture their dialogue, highlighting the playful and inquisitive spirit of their interaction.
@@
Response B
(Qwen2.5-VL-7B-Instruct)
This image is a creative and surreal depiction that combines elements of everyday objects with imaginative scenarios. It shows a large, tilted bottle pouring liquid onto the ground. The liquid spreads out in a pool around two women who are sitting on the edge of the puddle. One woman is partially submerged in the liquid, while the other sits on the edge looking at her. The scene is contrasted by the dark, plain background, which makes the bottle and the women stand out prominently. This type of imagery is often used in advertising or artistic representations to convey a message or evoke emotions.
@@
Response B
(Gemini-2.5-Flash)
Here’s a breakdown of the vehicles in the image: - There is one green bus in the lower right portion of the image. - There is a gray dirt bike in the middle-left. - There is a golden motorcycle (possibly a chopper or custom bike) with red and green accents in the upper-middle. - There is a gray airplane with blue and green accents in the upper-right. Based on this observation: - Number of green buses = 1 - Number of blue school buses = 0 (There are no blue buses, nor any vehicles that distinctly look like a school bus of any color, let alone blue.) Since 1 is greater than 0, the number of green buses is greater than the number of blue school buses. The final answer is $\boxed{A}$
@@ E.1 To Think or Not?
To study how thinking influences judge model behaviors in following diverse criteria, we examine two strong open-source LMM families that are known for producing high-quality reasoning traces across diverse domains. InternVL3.5 [wang2025internvl3.5] supports both thinking (default) and non-thinking modes controlled via system prompts, enabling a clean mode-switch comparison within the same model weights. Qwen3-VL [yang2025qwen3] provides two model variants—Instruct and Thinking—the latter specifically developed to enhance reasoning capabilities.
@@ E.2 Joint Multi-Criterion Judgment
    Joint multi-criterion judgment affects models unevenly. GPT-4o exhibits a clear performance drop across all metrics, suggesting that combining multiple criteria in a single pass amplifies its internal biases and weakens its ability to follow diverse criteria. GPT-5, in contrast, benefits from the joint setting—showing improved pluralistic accuracy and better alignment with human preferences on conflict cases. For o4-mini and Gemini-2.5-Pro, the effects of joint criterion judgment are mixed on open-ended tasks but consistently drop on reasoning domains.
@@
    Joint judgment generally reduces the model’s sensitivity to criterion-level trade-offs. Across all models—except GPT-5 in the reasoning split—the trade-off sensitivity decreases. This trend is expected: given the autoregressive nature of LLMs/LMMs, generating multiple criterion-level judgments within a single inference pass inevitably induces inter-criterion dependencies, making the judge model more likely to assign the same preference direction across criteria for the same pair of responses. Table 19 shows a case where o4-mini fails to capture criterion-level conflicts under joint criterion judgment.

### v2

@@ D.2 Full Evaluation Instances
Response B (yi-vl-plus) In a vividly colored comic book panel, we see two characters engaging in a lively conversation. One character, with a bright yellow head and a red shirt, exclaims, “An inventor? I don’t believe you!” His name, Marco, is displayed in a white rectangle. Opposite him, another character, sporting a black hairstyle and a blue shirt, confidently asks, “Why not?” His name, Babi, is also shown in a white rectangle. The contrasting colors of their clothes and hair add to the dynamic energy of their exchange, as they stand against a plain blue backdrop. The speech bubbles capture their dialogue, highlighting the playful and inquisitive spirit of their interaction.
@@
Response B (Qwen2.5-VL-7B-Instruct) This image is a creative and surreal depiction that combines elements of everyday objects with imaginative scenarios. It shows a large, tilted bottle pouring liquid onto the ground. The liquid spreads out in a pool around two women who are sitting on the edge of the puddle. One woman is partially submerged in the liquid, while the other sits on the edge looking at her. The scene is contrasted by the dark, plain background, which makes the bottle and the women stand out prominently. This type of imagery is often used in advertising or artistic representations to convey a message or evoke emotions.
@@
Response B (Gemini-2.5-Flash) Here’s a breakdown of the vehicles in the image: - There is one green bus in the lower right portion of the image. - There is a gray dirt bike in the middle-left. - There is a golden motorcycle (possibly a chopper or custom bike) with red and green accents in the upper-middle. - There is a gray airplane with blue and green accents in the upper-right. Based on this observation: - Number of green buses = 1 - Number of blue school buses = 0 (There are no blue buses, nor any vehicles that distinctly look like a school bus of any color, let alone blue.) Since 1 is greater than 0, the number of green buses is greater than the number of blue school buses. The final answer is $\boxed{A}$
@@ E.1 To Think or Not?
To study how thinking influences judge model behaviors in following diverse criteria, we examine two strong open-source LMM families that are known for producing high-quality reasoning traces across diverse domains. InternVL3.5 [43] supports both thinking (default) and non-thinking modes controlled via system prompts, enabling a clean mode-switch comparison within the same model weights. Qwen3-VL [55] provides two model variants—Instruct and Thinking—the latter specifically developed to enhance reasoning capabilities.
@@ E.2 Joint Multi-Criterion Judgment
    Joint multi-criterion judgment affects models unevenly.
    GPT-4o exhibits a clear performance drop across all metrics, suggesting that combining multiple criteria in a single pass amplifies its internal biases and weakens its ability to follow diverse criteria. GPT-5, in contrast, benefits from the joint setting—showing improved pluralistic accuracy and better alignment with human preferences on conflict cases. For o4-mini and Gemini-2.5-Pro, the effects of joint criterion judgment are mixed on open-ended tasks but consistently drop on reasoning domains.
@@
    Joint judgment generally reduces the model’s sensitivity to criterion-level trade-offs.
    Across all models—except GPT-5 in the reasoning split—the trade-off sensitivity decreases. This trend is expected: given the autoregressive nature of LLMs/LMMs, generating multiple criterion-level judgments within a single inference pass inevitably induces inter-criterion dependencies, making the judge model more likely to assign the same preference direction across criteria for the same pair of responses. Table 19 shows a case where o4-mini fails to capture criterion-level conflicts under joint criterion judgment.